OpenStudy (anonymous):

Determine which ordered pair in the form (x, y) is a solution of the given equation. (x + 3)y = 14 (Points : 1) (11, 1) (3, 2) (5, 2) (7, 2)

OpenStudy (anonymous):

Which pairs are solutions to the equation? 5xy + 9 = 44 (Points : 1) (1, 7) and (4, 11) (1, 7) and (7, 1) (7, 1) and (2, 3) (4, 11) and (2, 3) this is the last Q of the test

OpenStudy (anonymous):

so the first question: you have 4 choices of x and y, if you substitute those into the equation, it is a solution if it makes it true, such as 2x=2, a solution to this is x=1 because then 2=2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what about the second?

OpenStudy (anonymous):

so if you start with the first one put 11 in for x and 1 in for y. does (11+3)(1) = 14? if so then that is the answer

OpenStudy (anonymous):

do the same for the second. just now you have more numbers to try.

OpenStudy (anonymous):

ya it does

OpenStudy (anonymous):

for the second one. look at the first one 5xy + 9 = 44 x = 1 y = 7 and the pair x = 4 y = 11 5(1)(7) + 9 = 44 and 5(4)(11) + 9 = 44 5*7 + 9 = 44 and 5* 44 + 9 = 44 35+9 = 44 and 220+9 = 44 then you ask do both of these make sense? if not then go to the next set of pairs

OpenStudy (anonymous):

by first one here i mean the first ordered pairs

OpenStudy (anonymous):

to me it kinda makes sense does it make sense to you

OpenStudy (anonymous):

it doesnt because 220+9 =44 isnt true. 229 =44 and 229 does not equal 44. but you do get some help here though. after looking at those first pairs you found that x=1 and y = 7 works to have 44=44. so if you try others with that ordered pair you already know one of those works. looking at the second set x=1 y=7 and x=7 y=1

OpenStudy (anonymous):

man!donts you wish math and algebra and all kinds of math wasnt invented so hard

OpenStudy (anonymous):

haha i like math, i wish english wasnt here. i mean i like speaking and all but get rid of the nouns and verbs. english isnt my thing

OpenStudy (anonymous):

math isnt my thing AT ALL

OpenStudy (anonymous):

ok my test thing says i only have a minute to finsh or it will be turned in and i have tow Qs left but im not done ugh

OpenStudy (anonymous):

two**

OpenStudy (anonymous):

do you know the answer to the 2nd question you asked

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i have 30 seconds left

OpenStudy (anonymous):

okay so you know x=1 and y=7 works and since it is multiplication y=7 and x=1 works as well

OpenStudy (anonymous):

so b

OpenStudy (anonymous):

my math grade is low thats why i want some good correct answers

OpenStudy (anonymous):

and the first one was the first option

OpenStudy (anonymous):

ok thank you so much for helping me

OpenStudy (anonymous):

no problem